### v4

# EF-3DGS: Event-Aided Free-Trajectory 3D Gaussian Splatting

Abstract
Scene reconstruction from casually captured videos has wide real-world applications. Despite recent progress, existing methods relying on traditional cameras tend to fail in high-speed scenarios due to insufficient observations and inaccurate pose estimation. Event cameras, inspired by biological vision, record pixel-wise intensity changes asynchronously with high temporal resolution and low latency, providing valuable scene and motion information in blind inter-frame intervals. In this paper, we introduce the event cameras to aid scene construction from a casually captured video for the first time, and propose Event-Aided Free-Trajectory 3DGS, called EF-3DGS, which seamlessly integrates the advantages of event cameras into 3DGS through three key components. First, we leverage the Event Generation Model (EGM) to fuse events and frames, enabling continuous supervision between discrete frames. Second, we extract motion information through Contrast Maximization (CMax) of warped events, which calibrates camera poses and provides gradient-domain constraints for 3DGS. Third, to address the absence of color information in events, we combine photometric bundle adjustment (PBA) with a Fixed-GS training strategy that separates structure and color optimization, effectively ensuring color consistency across different views. We evaluate our method on the public Tanks and Temples benchmark and a newly collected real-world dataset, RealEv-DAVIS. Our method achieves up to 3dB higher PSNR and 40% lower Absolute Trajectory Error (ATE) compared to state-of-the-art methods under challenging high-speed scenarios.
Figure 1: Free-trajectory 3DGS under high speed. (Top) The overall paradigm. The colored dots in the top row represent the event data (red: positive, blue: negative). We leverage continuous event streams to aid discrete video frames captured along free trajectories in high-speed scenarios, jointly optimizing camera poses and reconstructing the 3DGS. Our method surpasses current state-of-the-art methods in terms of both rendered results (middle) and pose estimation (bottom).
1 Introduction
In recent years, Neural Radiance Fields (NeRF) nerf ; mipnerf360 ; instantngp and 3D Gaussian splatting (3DGS) 3dgs ; cf3dgs have made significant progress in novel view synthesis tasks. Given a set of posed images of the same scene, they optimize an implicit or explicit scene representation using volume rendering. While subsequent methods mipnerf360 ; instantngp ; tensorf ; zipnerf excel with posed images, reconstructing scenes from videos with free camera trajectories remains challenging despite its applications in VR/AR, video stabilization, and mapping. To tackle this challenging task, several efforts have been made.
Accurate pose estimation is often difficult to obtain in free-trajectory scenarios, which directly impacts the quality of scene reconstruction. One line of work draws inspiration from Simultaneous Localization and Mapping (SLAM). They localrf ; cf3dgs ; cogs follow its optimization paradigm, progressively optimizing camera trajectories and alternating between camera pose and scene refinement. Another line of work nopenerf ; cogs ; rodynrf ; cf3dgs ; sparf explores incorporating additional geometric or motion priors such as depth estimation depth1 ; depth2 or optical flow raft to establish constraints beyond photometric rendering loss. While these methods can render photo-realistic images in typical free-trajectory scenarios, both their rendering quality and pose estimation accuracy degrade significantly in high-speed scenarios (or equivalently low-frame-rate scenarios) as shown in Fig. 1. Such high-speed scenarios have essential applications such as autonomous driving and First-Person View (FPV) exploration.
The performance degradation of prior methods can be attributed to two primary factors. First, the limited number of camera observations leads to an under-constrained scene reconstruction problem. This can cause the scene representation to converge to a trivial solution regnerf ; sparsenerf ; sparf , where the model overfits to the training views without capturing the correct underlying geometry structure. Second, the substantial discrepancies between consecutive frames, resulting in diminished overlapping regions, violate the implicit assumption of continuous motion between adjacent frames, which is leveraged by previous methods. Moreover, geometric and motion priors like optical flow and feature matching become unreliable in such scenarios. These significant violations greatly exacerbate the ill-posedness of the joint optimization of scene and camera poses.
Event camera is a bio-inspired image sensor that asynchronously records per-pixel brightness changes, offering advantages such as high temporal resolution, high dynamic range, and no motion blur event_survey ; EDSO ; ultimateslam ; wan2024event ; wu2024event ; chengzhi . The brightness information recorded in the event stream can effectively complement the missing scene information between consecutive frames. Moreover, the event data naturally encodes the motion information of the scene EvMotionEstimation ; ganchao ; 10047993 , containing rich motion cues. These properties make event cameras well-suited for scene reconstruction tasks in high-speed and free-trajectory scenarios. However, seamlessly integrating the aforementioned benefits of event cameras is nontrivial. First, 3DGS renders absolute pixel brightness, which aligns with image data. Event cameras, however, record sparse differential brightness changes. Directly integrating the differential operations into 3DGS may amplify noise and lead to ill-conditioned optimization problems with high sensitivity to parameter initialization and perturbations. Second, event cameras encode motion through continuous spatio-temporal trajectories of events. In contrast, frame-based data inherently discretizes continuous motion, forcing traditional methods to rely on correspondence matching, which fails in high-speed scenarios with large inter-frame displacements. These fundamental challenges require carefully designed method that bridges the gap between the event data and 3DGS optimization.
In this work, we propose Event-Aided Free-Trajectory 3DGS, dubbed EF-3DGS, a framework that integrates event data into the scene optimization process to fully leverage its high temporal resolution property. Our approach comprises three key components: (1) In the Event Generation Model (EGM), we introduce an event-based re-render loss, which extends the 3DGS optimization to the continuous event stream. This allows us to utilize the brightness cues encoded in the event stream between adjacent frames, providing rich supervisory signals to alleviate the insufficient sparse view issues. (2) In the Linear Event Generation Model (LEGM), regarding the pose estimation challenge, we introduce the CMax cmax framework to exploit the spatio-temporal correlations of events. We obtain the motion field by leveraging the pseudo-depth from 3DGS rendering and the relative camera motion between consecutive frames. We then warp the events triggered by the same edge along the motion trajectories to maximize the sharpness of the image of warped events (IWE), thereby estimating the motion that best matches the current spatio-temporal event patterns. Furthermore, through the LEGM legm2015 ; legm2018 , we establish a connection between motion and brightness changes. This allows us to constrain the 3DGS in the gradient domain using the IWE. (3) As most event data primarily records scene brightness changes, lacking color information, we introduce photometric bundle adjustment (PBA) and a Fixed-GS strategy to address this. PBA recovers color by optimizing reprojection errors onto RGB frames, while Fixed-GS enables separate optimization of scene structure and color.
Our main contributions are summarized as follows:
  • We introduce event cameras into the task of free-trajectory scene reconstruction for the first time. Its advantage of high temporal resolution and low latency showcases the potential of event data for scene reconstruction tasks in challenging scenarios.
  • We derive our method from the underlying imaging principles of event cameras and design the corresponding loss functions that mine the motion and brightness information encoded in event data and seamlessly integrate them into the 3DGS optimization.
  • Experiments on both public benchmarks and real-world datasets demonstrate that our method significantly outperforms existing state-of-the-art approaches in terms of both rendering quality and trajectory estimation accuracy.
2 Related Works
Joint Pose and Scene Optimization. The research community has recently focused on developing methods sparf ; localrf ; nerfmm ; graf ; nopenerf ; barf ; cf3dgs that can be optimized without requiring precomputed camera poses. A line of work has focused on improving the stability of the optimization process. GARF graf and BARF barf both find that the high-frequency position encoding is prone to local minima and try to improve it. For example, GARF graf proposes using Gaussian activation to replace the sinusoidal position encoding. Another line of work has investigated incorporating additional constraints to make the problem more tractable. LocalRF localrf leverages the prior assumption of continuous motion between adjacent frames and progressively adds and optimizes camera poses. More recent approaches nopenerf ; localrf ; cf3dgs leverage pre-trained networks, i.e., monocular depth estimation and optical flow estimation. Exploiting 3DGS’s explicit representation, CF-3DGS cf3dgs directly back-projects Gaussian points using depth maps. While the aforementioned methods have made notable progress, they have yet to fully address the challenges posed by high-speed scenarios or rely on a good pose initialization. Our approach addresses these issues by leveraging motion and brightness cues from event streams.
Event-Based Novel View Synthesis. Recent works have explored the integration of event cameras EMoTive ; etap ; urnet into the NeRF or 3DGS framework. Early approaches, such as E-NeRF enerf and EventNeRF eventnerf , utilize event-based generative models, minimizing the difference between the rendered brightness changes and observed brightness changes. Building upon this, Robust e-NeRF robustenerf incorporates a more realistic imaging model into the event-based framework, accounting for factors like refractory periods and noise. Beyond event-based NeRF, efforts have also been made to integrate event data into image-based methods. For instance, E2NeRF E2NeRF and EvDeblurNeRF evdeblurnerf leverage the Event Double Integral (EDI) edi model to address the deblurring problem, while DE-NeRF denerf and EvDNeRF EvDNeRF leverage the high temporal resolution property of event cameras to capture fast-moving elements in dynamic scene. More recently, Event-3DGS event3dgs and EaDeblur-GS eadeblurgs have extended previous approaches to 3D Gaussian Splatting, achieving superior rendering quality and real-time performance. A key distinction of our work is that, unlike the prior methods that rely on accurate precomputed poses, we target free-trajectory scenarios, jointly optimizing for both the camera poses and the scene representation. Furthermore, while previous works have been limited to simulated and simple environments, we evaluate our approach in large-scale outdoor scenarios with complex motions and lighting conditions.
Figure 2: Method overview. The inputs are video frames and event stream. In the first stage, we progressively add new event images, leveraging the events and most recent frame to establish the event-driven optimization. In the second stage, we adopt the Fixed-GS strategy to mitigate the color distortion of 3DGS. The details of $\mathcal{L}_{LEGM}$ and CMax framework are shown in Fig. 3.
3 Preliminary
3DGS 3dgs parametrizes the 3D scene as a set of 3D gaussians $\{G_{k}\}_{k=1}^{K}$ that carry the geometric and appearance information. Each 3D Gaussian is characterized by several learnable properties, including its center position $\mu\in\mathbb{R}^{3}$, opacity $\alpha\in[0,1]$, spherical harmonics (SH) features $\mathbf{f}_{k}\in\mathbb{R}^{3\times 16}$ for view-dependent color $c\in\mathbb{R}^{3}$, rotation matrix $R\in\mathbb{R}^{3\times 3}$ (stored in quaternion form), scale factor $s\in\mathbb{R}^{3}$. The shape of each Gaussian is defined by the covariance matrix $\mathbf{\Sigma}$ and the center (mean) point $\mu$, $G(x)=\exp(-\frac{1}{2}{(x-\mu)}^{T}\mathbf{\Sigma}^{-1}(x-\mu)).$ During rendering, a tile-based rasterizer is applied to enable fast sorting and $\alpha$-blending. The color of each pixel is calculated via blending N ordered overlapping points:
$C(\textbf{r})=\sum_{i=1}^{N}c_{i}\alpha_{i}\ \prod_{j=1}^{i-1}(1-\alpha_{j}),$ (1)
where $c_{i}$ is calculated from spherical harmonics and view direction, $\alpha_{i}$ is the multiplication of opacity and the transformed 2D Gaussian and r denotes the image pixel. With the forward rendering procedure, we can optimize 3DGS by minimizing a weighted combination loss of $\mathcal{L}_{1}$ and $\mathcal{L}_{D-SSIM}$ between observation and rendered pixels: $\mathcal{L}_{color}=(1-\lambda)\mathcal{L}_{1}(\hat{I},I)+\lambda\mathcal{L}_{D-SSIM}(\hat{I},I),$ where $\lambda$ is balancing weight which is set to 0.2 following 3dgs . By integrating depth $d_{i}$ in Equation˜1 along the ray, we can also obtain a expected depth value $\hat{D}(\mathbf{r})$:
$\hat{D}(\mathbf{r})=\sum_{i=1}^{N}d_{i}\alpha_{i}\ \prod_{j=1}^{i-1}(1-\alpha_{j}).$ (2)
4 Method
[…]
4.1 EGM Driven Optimization
The EGM describes how event cameras asynchronously record pixel-wise brightness changes. When the logarithmic brightness change at a pixel $\mathbf{u}_{k}=(x_{k},y_{k})$, exceeds a predefined contrast threshold $C$,
$\Delta L(\mathbf{u}_{k},t_{k})\doteq L(\mathbf{u}_{k},t_{k})-L(\mathbf{u}_{k},t_{k}-\delta t)=p_{k}C,$ (3)
where $L\doteq log(I)$ is the logarithm of intensity, $p_{k}\in\{-1,+1\}$ indicates the polarity of brightness changes, and $t_{k}$ is the triggered timestamp.
As shown in Fig. 2 (1.1), to leverage the high temporal resolution of events, we first divide the time interval between two adjacent video frames $I_{i}$ and $I_{i+1}$ into N smaller subintervals $\varepsilon_{i,j}=\{\mathbf{e}_{k}|t_{i,j}\leq t_{k}\leq t_{i,j+1},\Delta t=\frac{t_{i+1}-t_{i}}{N},t_{i,j}=t_{i}+j\cdot\Delta t\}$. This allows us to form accumulated event frames at a higher temporal resolution:
$E_{i,j}=\sum_{e_{k}\in\varepsilon_{i,j}}p_{k}.$ (4)
We then reconstruct the latent intensity image $I_{t}$ at any intermediate time $t\in\{t_{i,j}\}$ by integrating the accumulated events with the most recent frame:
$I_{t}=I_{i,j}=\begin{cases}I_{i,0}\cdot\exp(\sum_{n=0}^{j-1}E_{i,n}\cdot C)&\text{if }j>0\\ I_{i,0}&\text{if }j=0\end{cases}.$ (5)
This latent intensity image provides a supervisory signal for our event-based rendering loss:
$\mathcal{L}_{EGM}=(1-\lambda)\mathcal{L}_{1}(\hat{I}_{t},I_{t})+\lambda\mathcal{L}_{D-SSIM}(\hat{I}_{t},I_{t}).$ (6)
By enforcing consistency between rendered and latent intensity images, this loss effectively utilizes the brightness information encoded in event streams between adjacent frames, addressing the challenge of sparse viewpoints in high-speed scenarios.
Figure 3: The illustration of unified CMax and LEGM optimization. We warp previous event frames to the sampled timestamp through the optical flow and maximize the sharpness of the image of IPWE. The byproduct IPWE is utilized to establish additional constraints on 3DGS.
4.2 Unified CMax and LEGM Optimization
While $\mathcal{L}_{EGM}$ leverages the brightness change information recorded by events, it does not explicitly exploit the motion information encoded in the event stream. To address this, we introduce the Contrast Maximization (CMax) cmax ; cmax2 ; cmax3 ; cmaxslam framework and the LEGM legm2015 ; legm2018 ; linearinversed . These models complement the previous EGM-driven optimization.
Under constant scene illumination, events are triggered by the motion of scene edges, forming continuous trajectories in (x, y, t) space. As shown in Fig. 2 (1.2), by warping(back-projecting) events along the correct motion trajectories, we can obtain a sharp image of warped events (IWE). Therefore, the sharpness of the IWE can serve as an indication of the accuracy of the estimated motion. This insight motivates us to derive the motion field by leveraging the rendered depth from 3DGS using eq.˜2 and the relative camera motion between neighboring timestamps. By optimizing the sharpness of the IWE, we can obtain the optimal motion field, which in turn helps to improve the geometric accuracy of the 3DGS and the camera poses.
As shown in Fig. 3, for efficiency, we adopt a piece-wise warping approach instead of warping individual events. Specifically, for current timestamp $t_{ref}=t_{i,j}$, we warp the event frames from previous $r$ sub-intervals:
$E_{i,j-m\rightarrow j}=\textrm{warp}(E_{i,j-m},F_{i,j\rightarrow j-m}),$ (7)
where $m\in[0,r]$, $F_{i,j-m\rightarrow j}$ is the optical flow derived from the rendered depth $\hat{D}$ in  eq.˜2 and relative pose $T_{i,j\rightarrow j-m}$ between two timestamps:
$F_{i,j\rightarrow j-m}=\Pi(T_{i,j\rightarrow j-m}\Pi^{-1}(x,y,\hat{D}))-(x,y),$ (8)
where $T_{i,j\rightarrow j-m}=T_{i,j-m}T_{i,j}^{-1}$, $\Pi$ projects a 3D point to image coordinates and $\Pi^{-1}$ unprojects a pixel coordinate and depth into a 3D point. Then the image of piece-wise warped events (IPWE) at timestamp $t_{i,j}$ is computed by averaging the warped event frames:
$\textrm{IPWE}_{i,j}=\frac{1}{r+1}\sum_{m=j-r}^{j}E_{i,m\rightarrow j}\approx\frac{1}{C}\Delta L.$ (9)
Following the Cmax framework, we maximize the variance of the IPWE, which is equivalent to minimize its opposite:
$\mathcal{L}_{cm}=-\textrm{Var}(\textrm{IPWE}_{i,j}).$ (10)
Furthermore, based on the LEGM legm2015 ; legm2018 , the brightness change $\Delta L$ at pixel $\mathbf{u}$ can be approximated by the dot product of the image gradient $\nabla L$ and the optical flow $\dot{\mathbf{u}}$ (note that $L$ is the logarithm of an image):
$\Delta L(\mathbf{u})=-\nabla L\cdot\dot{\mathbf{u}}\approx L(\mathbf{u})-L(\mathbf{u}+\dot{\mathbf{u}}).$ (11)
It is noteworthy that the IPWE also encodes brightness change information. Combining  eq.˜9 and  eq.˜11, we establish a connection between the IPWE and the brightness changes of the rendered images:
$C\cdot\textrm{IPWE}_{i,j}=\hat{L}(\mathbf{u})-\hat{L}(u+F_{i,j\rightarrow j+1}).$ (12)
Note that to compute $F_{i,j\rightarrow j+1}$, we estimate $T_{i,j+1}$ by leveraging the assumption of locally linear motion from $T_{i,j-1}$ and $T_{i,j}$. Based on this relationship, we formulate an additional gradient-based loss:
[…]
where $\hat{L}$ is the logarithm of synthesised image $\hat{I}_{t}$. Finally, the full LEGM loss is defined as:
$\mathcal{L}_{LEGM}=\lambda_{cm}\mathcal{L}_{cm}+\lambda_{grad}\mathcal{L}_{grad},$ (14)
where $\lambda_{cm}$ and $\lambda_{grad}$ are the balancing weight.
4.3 Photometric Bundle Adjustment
The aforementioned event-based constraints, $\mathcal{L}_{EGM}$ and $\mathcal{L}_{LEGM}$, leverage the brightness change and motion information encoded in the event data to constrain 3DGS. However, as event cameras only record brightness changes and lack color perception, directly applying them to 3DGS optimization may lead to inconsistent color rendering. To ensure cross-view consistency of the 3DGS rendering, we introduce the Photometric Bundle Adjustment (PBA) term.
Specifically, as shown in Fig. 2 (1.3), for a randomly sampled timestamp $t\in\{t_{i,j}\}$, we establish the following photometric reprojection error:
[…]
where $\mathbf{u}^{\prime}=\Pi(T_{i,j-r\rightarrow j}\Pi^{-1}(x,y,\hat{D}(u)))$ represent the coordinate on target view projected from the pixel $\mathbf{u}$ of source view $I_{s}$, $\mathcal{P}$ denotes the pixel samples of current frame, and $\mathcal{F}$ is the candidates of target video frames. We select $\mathcal{F}$ to be the nearest previous video frame in consideration of computation costs.
By minimizing $\mathcal{L}_{PBA}$ across sampled views, we encourage the 3DGS model to produce geometrically and photometrically consistent renderings across events and video frames, thus effectively resolving color inconsistencies inherent in event data.
4.4 Fixed-GS Training Strategy
The $\mathcal{L}_{PBA}$ term alone is insufficient to fully mitigate color distortion issues. To further address this challenge, we propose a two-stage Fixed-GS scene optimization strategy that takes advantage of 3DGS’s explicit attribute representation. In the first stage, all the parameters are optimizable and the optimization is performed across all timestamps:
$G_{\theta}^{*},T_{i,j}^{*}=\underset{\mu,\alpha,r,s,f,T_{i,j}}{\operatorname{argmin}}~\mathcal{L}_{event},~t\in\{t_{i,j}\},$ (16)
where $\mu,\alpha,r,s,f$ is the position, opacity, rotation, scale factor and spherical harmonics of the Gaussians, and t is the sampled timestamp during training. This stage results in a scene reconstruction with accurate structure and brightness, albeit with potential color distortions due to the dominant colorless event supervision overwhelming the sparse RGB frame color supervision. The second stage focuses on recovering accurate color information. During this phase, optimization is conducted exclusively on video frames. We optimize only the spherical harmonic coefficients of the Gaussians while keeping other parameters fixed:
$G_{\theta}^{*}=\underset{f}{\operatorname{argmin}}~\mathcal{L}_{color},~t\in\{t_{i,0}\}$ (17)
The ratio between the first and second stages is empirically set to 4:1. This approach allows us to effectively address the color distortion problem while preserving the structural and brightness information obtained from the event data.
4.5 Overall Training Pipeline
Assembling all loss terms, we get the overall loss function:
$\mathcal{L}_{event}=\mathcal{L}_{EGM}+\mathcal{L}_{LEGM}+\lambda_{PBA}\mathcal{L}_{PBA},$ (18)
where $\lambda_{PBA}$ are the weighting factor. Note that since event cameras typically record only the changes in brightness intensity, the $\mathcal{L}_{EGM}$ and $\mathcal{L}_{LEGM}$ losses are computed in the grayscale domain, whereas the $\mathcal{L}_{PBA}$ loss is calculated in RGB color space. We incorporate dynamic scene allocation strategies from LocalRF localrf for handling extended video sequences. Our overall training pipeline builds upon the progressive optimization scheme of CF-3DGS cf3dgs while introducing novel components to integrate event stream data for robust free-trajectory scene reconstruction. Please refer to Section˜A.3 for the algorithm pipeline and additional implementation details.
5 Experiments
Methods Pose-Free Input 6 FPS 4 FPS 3 FPS 2 FPS 1 FPS
[…]
F2-NeRF $\times$ F 23.55 0.75 0.34 22.97 0.72 0.36 22.25 0.69 0.40 21.64 0.68 0.44 20.63 0.64 0.51
[…]
EvDeblurNeRF $\times$ E+F 22.43 0.71 0.38 21.23 0.69 0.42 20.09 0.65 0.49 17.52 0.62 0.55 15.19 0.55 0.60
ENeRF $\times$ E+F 23.62 0.73 0.37 22.84 0.70 0.38 21.85 0.69 0.41 20.52 0.66 0.46 18.09 0.60 0.52
Event-3DGS(E+F) $\times$ E+F 26.32 0.78 0.33 25.37 0.76 0.34 24.59 0.75 0.37 23.44 0.72 0.38 22.41 0.69 0.39
[…]
Table 1: Quantitative evaluations on Tanks and Temples dataset. The best results are highlighted in bold. Note that, in the “input” column, “F” denotes traditional frame input, while “E+F” denotes hybrid frame and event input.
Methods Input 6 FPS 4 FPS 3 FPS 2 FPS 1 FPS
[…]
Nope-NeRF F 0.1141 0.7563 2.8382 0.1604 1.0542 2.7653 0.2220 1.3694 2.7857 0.3131 1.8965 2.8412 0.6216 3.913 2.6592
LocalRF F 0.0806 0.9282 0.5630 0.0867 0.9683 0.6085 0.0911 0.9800 0.6501 0.0957 1.0428 0.6802 0.1421 1.4725 1.0006
CF-3DGS F 0.0594 0.6981 0.4212 0.0637 0.7128 0.4628 0.0712 0.7531 0.5189 0.0859 0.8074 0.6918 0.1057 0.9768 0.8972
EvCF-3DGS E+F 0.0461 0.5972 0.3419 0.0490 0.6269 0.3766 0.0538 0.6728 0.4261 0.0591 0.7094 0.4860 0.0657 0.7597 0.5534
EF-3DGS(Ours) E+F 0.0391 0.5427 0.2885 0.0407 0.5521 0.3064 0.0426 0.5796 0.3271 0.0449 0.5953 0.3671 0.0487 0.6259 0.3753
Table 2: Pose accuracy on Tanks and Temples. We use COLMAP poses in Tanks and Temples as the “ground truth”. The unit of $\textrm{RPE}_{r}$ is in degrees, ATE is in the ground truth scale and $\textrm{RPE}_{t}$ is scaled by 100. Those methods that require precomputed poses are excluded.
5.1 Dataset
Tanks and Temples.  We conduct comprehensive experiments on the Tanks and Temples dataset tnt . Similar to LocalRF localrf , we adopt 9 scenes, covering large-scale indoor and outdoor scenes. For each scene, we sample a video clip with a 50-second duration, typically featuring free camera trajectories and covering a considerable distance. Following LocalRF localrf , we apply 4× spatial downsampling to the videos. To evaluate the robustness under varying camera speeds, we employ varying temporal downsampling of 6 FPS, 4 FPS, 3 FPS, 2 FPS, and 1 FPS. The reduction in frame rate effectively creates larger inter-frame displacements, simulating high-speed scenarios. To synthesize realistic event data, we first upsample the original videos by RIFE and then apply the simulator V2E v2e .
RealEv-DAVIS.  Due to the lack of free-trajectory event camera datasets, we introduce RealEv-DAVIS. Using a DAVIS346 camera that simultaneously captures frames and events at 346×260 resolution, we record 40-second handheld sequences at 25 FPS. We employ COLMAP for ground-truth poses. For SLOW scenarios, we retain every second frame, while for FAST scenarios, we keep only one frame per five frames. Further details are provided in Section˜A.2.
5.2 Implementation details
We follow the optimization parameters by the configuration outlined in the 3DGS 3dgs . We optimize the camera poses in the representation of quaternion rotation. The initial learning rate is set to $10^{-5}$ and gradually decays to $10^{-6}$ until convergence. The balancing weight $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$ is empirically set to 0.1, 0.2 and 0.5. For the division of events between adjacent frames, we maintain a constant interval of $\frac{1}{6}$s for Tanks and Temples and $\frac{1}{25}$s for RealEv-DAVIS, setting the number of subinterval N accordingly. For example, in Tanks and Temples, N equals 2 for 3FPS and 6 for 1FPS. This ensures adherence to the constant brightness assumption within each sub-interval and provides adequate events for the following CMax warping. The intervals of neighboring warping $r$ in CMax are set to 3. The contrast threshold $C$ is set to 0.25 for Tanks and Temples and 0.21 for RealEv-DAVIS. We provide detailed ablation studies on these hyperparameters and additional implementation details in Section˜A.4.
Figure 4: Qualitative comparison for novel view synthesis. The first two rows are from Tanks and Temples and the last row is from RealEv-DAVIS. Our approach produces more realistic rendering results with fine-grained details. Better viewed when zoomed in.
Figure 5: Pose estimation comparison. We visualise the trajectory (3D plot) and $\mathrm{RPE}_{r}$ (color bar) of each method. We clip and normalize the $\mathrm{RPE}_{r}$ by a quarter of the max $\mathrm{RPE}_{r}$ across all results of each scene.
5.3 Experimental Setup
Metrics.  We evaluate all the methods from two aspects: novel view synthesis and pose estimation. For the novel view synthesis task, we report the standard metrics PSNR, SSIM ssim , and LPIPS lpips . For the pose estimation task, we adopt the Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) metrics pose_eval1 ; pose_eval2 , as delineated in nopenerf . Since these metrics are inherently influenced by frame rate, we upsample all estimated poses to a consistent temporal resolution before evaluation for fair comparison across different frame rate settings.
Baselines.  For a fair comparison, we focus on two categories of methods: (1) For frame-based approaches, we selected methods specifically addressing free-trajectory scenarios, such as LocalRF localrf and F2-NeRF f2nerf . We also include pose-free methods like Nope-NeRF nopenerf and CF-3DGS cf3dgs . (2) For event-frame hybrid methods, we consider approaches that fuse events and frames, including ENeRF enerf , EvDeblurNeRF evdeblurnerf and Event-3DGS event3dgs .
Methods Input SLOW FAST
NVS Pose NVS Pose
[…]
LocalRF F 20.83 0.6074 3.60 2.07 17.62 0.5192 5.22 2.96
CF-3DGS F 22.68 0.6287 2.49 1.55 17.59 0.5204 3.68 2.17
EvDeblurNeRF E+F 20.61 0.6064 - - 17.98 0.5269 - -
Event-3DGS (E+F) E+F 23.43 0.6456 - - 20.04 0.5515 - -
EvCF-3DGS E+F 22.89 0.6317 1.78 0.82 19.13 0.5380 2.70 1.28
EF-3DGS(Ours) E+F 23.65 0.6466 1.41 0.69 21.12 0.5620 1.80 0.89
Table 3: Rendering and pose estimation results on RealEv-DAVIS. Complete data and additional metrics are provided in the supplementary material.
Since no existing method integrates events for free-trajectory scenarios, we implement EvCF-3DGS as a competitive baseline that leverages an event-based frame interpolation network (Time Lens timelens ) to temporally upsample frames before feeding them into CF-3DGS.
5.4 Experimental Results
We select every ten frames as a test image for NVS evaluation following LocalRF localrf . Since the camera poses are unknown in our setting, we need to estimate the poses of test views. As in iNeRF iNeRF , we freeze the 3DGS model, initialize the test poses with the poses of the nearest training frames, and optimize the test poses by minimizing the photometric error between rendered images and test views.
Results on RealEv-DAVIS. Table 3 validates our approach on the real-world RealEv-DAVIS dataset. EF-3DGS outperforms top-performing methods and handles real-world scenes effectively. In FAST scenarios, our method shows nearly 1dB PSNR improvement over the best baselines. This confirms our advantage in high-speed scenarios where frame-based methods struggle. Fig. 4 and Fig. 5 show our method preserves fine details and maintains accurate trajectories even during rapid motion, addressing key limitations of traditional approaches.
$\mathcal{L}_{EGM}$ $\mathcal{L}_{LEGM}$ $\mathcal{L}_{PBA}$ Fixed GS NVS Pose
[…]
20.53 0.65 0.46 0.1057 0.9768 0.8972
22.16 0.68 0.42 0.0651 0.7529 0.5779
21.07 0.67 0.44 0.0830 0.8869 0.7231
20.96 0.65 0.46 0.0938 0.9875 0.9112
22.83 0.68 0.40 0.0523 0.6387 0.3981
23.46 0.70 0.37 0.0523 0.6387 0.3981
23.09 0.70 0.38 0.0487 0.6259 0.3753
23.96 0.72 0.36 0.0487 0.6259 0.3753
Table 4: Effect of each component in EF-3DGS. The best results are highlighted in bold.
Figure 6: Robustness of different methods to pose disturbance.
Results on Tanks and Temples. Tables 1 and 2 demonstrate two key findings: (1) Our event-aided approach achieves up to 3dB higher PSNR and nearly 40% lower trajectory error at 1FPS compared to frame-based methods, indicating the critical value of event data in high-speed scenarios. (2) Our method maintains 1.55dB PSNR advantage over Event-3DGS at 1FPS, confirming that our integration framework effectively exploits the nature of event data beyond merely using it. Fig. 4 shows our method produces sharper edges and finer textures, while Fig. 5 illustrates we achieve more accurate trajectory estimation.
Performance under Varying Camera Speeds  As shown in Table 2 and Table 3, while all methods degrade as the frame rate decreases, our approach shows remarkable resilience. The performance gap widens significantly at lower frame rates, with our PSNR advantage over CF-3DGS cf3dgs increasing from 0.61dB at 6FPS to 3.43dB at 1FPS. Notably, our method also consistently outperforms other event-based methods (EvCF-3DGS and Event-3DGS). This confirms not only the value of event data in challenging scenarios but also the superiority of our integration approach.
5.5 Ablation Studies
Effect of Each Component  Table 4 presents a comprehensive ablation study of our key components under the challenging 1FPS setting on Tanks and Temples. $\mathcal{L}_{EGM}$ serves as the foundation of our approach, providing substantial improvements in both rendering quality (+1.63dB PSNR) and pose accuracy by enabling rich supervision between discrete frames. Building upon this, $\mathcal{L}_{LEGM}$ extracts motion information from events and constrains 3DGS in the gradient domain, significantly improving pose estimation while modestly enhancing rendering quality. $\mathcal{L}_{PBA}$, though designed to address color inconsistency issues, not only improves rendering quality but also enhances pose estimation accuracy by establishing geometric and photometric consistency across views. The Fixed-GS training strategy, while having no impact on pose optimization, significantly improves rendering quality by effectively separating structure and color optimization. We provide more intuitive ablation visualizations in Section˜A.5.
Robustness to Pose Disturbance  To validate the robustness of different methods under inaccurate pose initialization, a common challenge in practical scenarios, we introduce varying degrees of perturbations to the initial camera poses estimated by COLMAP. Specifically, following BARF barf , we parametrize the camera poses $\mathbf{p}$ with the $\mathfrak{se}(3)$ Lie algebra. For each scene, we synthetically perturb the camera poses with additive noise $\delta\mathbf{p}\sim\mathcal{N}(\mathbf{0},n\mathbf{\textrm{I}})$, where n is the noise level. Then, each method is initialized with the noised poses, after which the optimization is performed. The results are illustrated in Fig. 6. Notably, Event-3DGS event3dgs , which lacks the capability to optimize camera poses, exhibits a drastic performance degradation as the magnitude of pose disturbances increases. This observation validates the critical importance of joint pose-scene optimization. Furthermore, Our proposed framework demonstrates superior tolerance across all perturbation levels. Even under significant noise, our method experiences substantially less degradation in both rendering quality and trajectory accuracy.
6 Conclusions
In this work, we propose Event-Aided Free-Trajectory 3DGS (EF-3DGS), a novel framework that seamlessly integrates event camera data into the task of reconstructing 3DGS from casually captured free-trajectory videos. Our method effectively leverages the high temporal resolution and motion information encoded in event streams to enhance the 3DGS optimization process, leading to improved rendering quality and accurate camera pose estimation. By introducing the Event Generation Model and Linear Event Generation Model, we bridge the gap between differential event data and absolute brightness rendering in 3DGS. The proposed photometric bundle adjustment and Fixed-GS strategy further ensure accurate color recovery and scene structure optimization. Extensive experiments on both public benchmarks and real-world datasets validate the effectiveness of our approach, demonstrating significant improvements over state-of-the-art methods in both rendering quality and trajectory estimation accuracy. Future work would explore self-adaptive parameter adjustment strategies to enhance the method’s versatility and ease of use across various reconstruction tasks.
7 Acknowledgements
This work is supported by the National Natural Science Foundation of China (NSFC) under Grants 62225207, 62436008 and 62306295. The AI-driven experiments, simulations and model training were performed on the robotic AI-Scientist platform of Chinese Academy of Sciences.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction accurately reflect the main contributions. Please refer to the main body of the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Justification: Please refer to Appendix A.9 for our discussions on the limitations.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide the necessary formulas and related explanations in the main text. Please see refer to Method section about the blur model in main paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Justification: We identify our baseline method and explain the implementation details where we differ from the baseline method in Experiment Setup 5.3.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: The datasets and baseline methods in this paper are publicly available. Our source code and the self-collect dataset will be openly accessible after the paper is accepted.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have provided necessary implementation details in Section˜5.3.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We follow the experimental settings of previous works to conduct experiments. Therefore, the comparison with previous works is fair.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We have provided the information of our computer resources in Section˜5.2 and Section˜A.4.3.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We ensure that our experiments comply with the NeurIPS Code of Ethics in all respects.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Please refer to our discussion on broader impacts in Section˜A.10.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks for misuse of released data or models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have appropriately cited related work and abide by their licenses and terms. Please refer to Section˜A.11.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Justification: We detail the data acquisition in Appendix A.2.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLM is only used for writing.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Appendix
A.1 Discussion on Motion Blur
Image blur arises from brightness integration during exposure time and occasionally occurs in low-light conditions or high-speed motion scenarios. While our method does not explicitly target image blur, this does not diminish its core contributions. We can draw an analogy to 2D vision tasks: image deblurring addresses blur-induced degradation, while video frame interpolation tackles discontinuity from excessive inter-frame motion. These represent distinct challenges in computer vision. Similarly, in 3D reconstruction, blur-related degradation and the challenges of sparse viewpoints with pose estimation constitute separate problem domains that often co-occur in high-speed scenarios but require distinct technical solutions. Our work specifically addresses the latter—fundamental issues that persist regardless of blur conditions and represent critical bottlenecks in high-speed 3D reconstruction.
Previous event-based scene reconstruction methods have primarily focused on blur reconstruction while overlooking the challenges of sparse viewpoints and inaccurate pose estimation. Our approach targets this gap by leveraging a novel fusion strategy to enhance traditional image-based 3DGS reconstruction, addressing fundamental limitations that affect reconstruction quality independent of blur artifacts.
For completeness, we conducted supplementary experiments to evaluate our method’s performance under motion blur conditions. We extend our method to handle motion blur by incorporating the Event Double Integration (EDI) model edi , which reconstructs sharp intensity from event data. Specifically, we reformulate the intensity term $I_{i,0}$ in Eq. (5) using the EDI formulation:
$\hat{I}_{i,0}=\frac{(2n+1)\cdot B_{i}}{\sum_{k=-n}^{n}\exp\left(C\cdot\sum_{z=0}^{k}E_{i,z}\right)},$ (19)
where $B_{i}$ represents the blurred intensity, $E_{i,z}$ denotes the accumulated events, $C$ is the contrast threshold, and $n$ defines the temporal integration window of blur averaging.
Experimental Setup: Experiments were performed on multiple scenes from the Tanks and Temples tnt with a frame rate of 2FPS. Motion-blurred frames were synthesized adopting the blur generation protocol from the GoPro-Blur dataset through gamma correction and multi-frame averaging operations. We average every 30 frames to simulate blurring. We selected EvDeblurNeRF evdeblurnerf , currently the best-performing open-source event-based deblur scene reconstruction method, as our comparison.
Results: As demonstrated in Tab. 5, our method achieves comparable reconstruction quality to EvDeblurNeRF while simultaneously performing pose estimation. Visual comparisons in Fig. 7 show that our method renders sharper novel views under motion blur conditions. Note that EvDeblurNeRF requires pre-computed COLMAP poses. This limitation significantly restricts practical applicability in high-speed scenarios where accurate pose estimation is challenging. These results demonstrate that our method maintains robust performance under blur conditions while addressing the fundamental pose estimation and sparse-view challenge.
Table 5: Comparison of different methods on pose estimation and rendering quality.
[…]
EvDeblurNeRF CF-3DGS 21.17 0.69 0.105 0.897
EvDeblurNeRF COLMAP 22.58 0.70 - -
EDI + EF-3DGS (Ours) EDI + EF-3DGS (Ours) 23.18 0.72 0.062 0.642
Figure 7: Qualitative comparison on motion deblur scene reconstruction. Our EDI + EF-3DGS method produces sharper results comparable to EvDeblurNeRF
A.2 Dataset
For the synthetic dataset, following LocalRF localrf , we choose nine static scenes from the Tanks and Temples dataset, which cover both indoor and outdoor environments. To construct a real-world dataset, we utilized a handheld DAVIS346 event camera to capture a series of extended video sequences, simulating free camera trajectories. Particularly for the scenes named building, we deliberately introduced significant camera motion during acquisition to emulate realistic camera shake scenarios. Details about these sequences are illustrated in Table 6, where Mean. rotation represents the mean relative rotation angle between two adjacent frames and Max. rotation denotes the maximum relative rotation angle between two frames in a sequence. We select a 50-second segment from each sequence to highlight scenarios with free-trajectory camera movements.
Table 6: Details of selected sequences on Tanks and Temples tnt and RealEv-DAVIS.
Scenes Seq. Length Mean. rotation (deg) Max. rotation (deg)
Tanks and Temples
Auditorium 300 2.91 54.55
Ballroom 300 5.73 179.81
Caterpillar 300 3.18 102.37
Church 300 2.44 37.22
Courtroom 300 8.40 177.96
M60 300 5.29 179.99
Museum 300 6.05 176.98
Panther 300 4.33 124.63
Train 300 4.80 108.18
RealEv-DAVIS
building 500 5.65 67.65
hall 500 3.88 106.73
corner 500 3.04 96.11
outdoor 500 2.36 77.18
A.3 Overall Training Pipeline
We detail the comprehensive training pipeline in Algorithm 1. Note that for notational simplicity, we consolidate the subscripts $(i,j)$ into a single index $k$. Our approach incorporates the dynamic radiance field allocation strategy from LocalRF localrf , which assigns a new radiance field when the current pose exceeds a distance threshold of 1 from the existing field. However, in practical implementations, we encountered the OOM(out of memory) problem with this approach. To address this, we adopted a modified strategy of allocating a new 3DGS every 50 frames, based on the sequence length. This adjustment ensures efficient memory utilization while maintaining the benefits of dynamic allocation. For pose optimization and scene reconstruction, we adopt strategies from CF-3DGS. And we incorporate event-driven brightness and motion coherence constraints, enabling robust reconstruction in challenging high-speed scenarios.
Algorithm 1 Overall Training Pipeline.
1:Input: event frames $\{I_{k}\}^{K}_{k=0}(k=i\cdot N+j)$.
2:Output: camera poses $\{T_{k}\}^{K}_{k=1}$, 3DGS $\{G_{\theta_{n}}\}$.
3:Initialize current center frame of 3DGS, $l=0$;
[…]
5:  $G_{\theta_{n}}$ $\leftarrow$ Initialize($I_{l}$) $\triangleright$ Following cf3dgs , initialize global 3DGS using the current center frame $I_{l}$
6:  for k in $[l+1,min(l+50,K)]$ do $\triangleright$ Optimize the 3DGS and the poses of next 50 frames
7:
8:   # Pose Estimation Step
9:   $G_{local}$ $\leftarrow$ Initialize($I_{k-1}$) $\triangleright$ Following cf3dgs , initialize local 3DGS using the previous frame $I_{k-1}$
10:   $T_{k}$ $\leftarrow$ Optimize($\mathcal{L}_{event}$, $T_{k}$) $\triangleright$ estimate camera pose $T_{k}$ by minimizing $\mathcal{L}_{event}$ at $T_{k}$
11:  end for
12:
13:  # Scene Reconstruction Step
14:  $\{G_{\theta_{n}}\}$ $\leftarrow$ Optimize($\mathcal{L}_{event}$, $T_{[l,l+50]}$) $\triangleright$ Optimizing 3DGS $G_{\theta_{n}}$ by minimizing $\mathcal{L}_{event}$ at $T_{[l,l+50]}$
15:  SHs of $\{G_{\theta_{n}}\}$ $\leftarrow$ Optimize($\mathcal{L}_{color}$, $T_{[l,l+50]}$)   $\triangleright$ Fixed-GS stage: optimizing SHs of 3DGS $G_{\theta_{n}}$ by minimizing $\mathcal{L}_{color}$ at RGB frames
16:  $l\leftarrow k$ $\triangleright$ Slide the current center frame to the next center frame
17:end while
Table 7: Ablation study results of parameter $r$ in CMax framework
   $r$ NVS Pose
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ $\text{RPE}_{t}\downarrow$ $\text{RPE}_{r}\downarrow$ ATE $\downarrow$
1 22.94 0.70 0.39 0.0591 0.6829 0.4851
3 23.96 0.72 0.36 0.0487 0.6259 0.3753
5 23.49 0.72 0.37 0.0524 0.6431 0.4338
A.4 Additional Experiments
A.4.1 Detailed Experiment Results on RealEv-DAVIS
Due to space constraints in the main text, we provide a detailed table of our method’s performance on the RealEv-DAVIS dataset in the supplementary materials, as shown in Tab. 8 and Tab. 9.
Table 8: Quantitative Evaluations on RealEv-DAVIS. We select the top-performing methods from the previous evaluation. Each baseline method is trained with its public code under the original settings and evaluated with the same evaluation protocol. The best results are highlighted in bold.
Methods Pose-Free Input SLOW FAST
hall building corner outdoor hall building corner outdoor
[…]
EvDeblurNeRF $\times$ E+F 20.24 0.5762 0.55 19.35 0.5649 0.50 20.47 0.5757 0.44 22.36 0.7087 0.38 18.23 0.5043 0.65 16.41 0.4624 0.63 17.82 0.4959 0.59 19.44 0.6448 0.42
ENeRF $\times$ E+F 21.89 0.5933 0.52 21.94 0.5918 0.36 22.01 0.5918 0.42 23.91 0.7266 0.30 20.09 0.5261 0.62 18.99 0.4903 0.50 19.61 0.5173 0.52 20.00 0.6494 0.42
Event-3DGS $\times$ E+F 23.01 0.6085 0.47 22.45 0.6191 0.34 23.11 0.6148 0.41 25.14 0.7399 0.23 20.35 0.5328 0.62 18.52 0.4907 0.51 20.05 0.5231 0.51 21.24 0.6596 0.40
[…]
Table 9: Pose accuracy on RealEv-DAVIS. The unit of $\textrm{RPE}_{r}$ is in degrees, ATE is in the ground truth scale and $\textrm{RPE}_{t}$ is scaled by 100. Those methods which require precomputed poses are excluded.
Methods Input SLOW FAST
hall building corner outdoor hall building corner outdoor
[…]
LocalRF F 4.4017 2.5067 0.2705 4.6186 1.9409 0.4909 1.9296 2.5183 0.4177 3.4307 1.3005 0.1984 5.6061 3.2265 0.3250 7.0611 2.8722 0.6853 3.4767 3.8950 0.5693 4.7229 1.8595 0.2798
CF-3DGS F 2.7106 0.6035 0.1459 3.0374 3.2761 0.2142 1.9437 1.4954 0.2225 2.2747 0.8053 0.0865 3.5025 0.7639 0.2635 4.6758 4.5256 0.4640 3.4950 2.2855 0.3482 3.0551 1.1198 0.1909
Ev-Baseline E+F 0.7063 0.4386 0.1218 3.4225 1.1717 0.1988 2.1283 1.3361 0.1565 0.8647 0.3455 0.0857 0.8898 0.7734 0.1307 5.0707 1.6997 0.3166 3.6785 2.1552 0.2121 1.1645 0.5001 0.1556
EF-3DGS(Ours) E+F 0.5262 0.3251 0.1041 2.8008 0.4633 0.1737 1.8386 1.6168 0.1864 0.4726 0.3699 0.0749 0.5424 0.3464 0.1369 3.5848 0.5589 0.2645 2.5396 2.2157 0.2029 0.5218 0.4462 0.1136
A.4.2 Evaluating the Influence of parameter r in LEGM
We investigate the impact of the parameter $r$ in the Contrast Maximization (CMax) framework, which determines the number of previous event frames warped to the current sampled timestamp. Table 7 presents the results of this ablation study. The optimal value for $r$ is found to be 3, yielding the best performance across all metrics for both novel view synthesis and pose estimation. When $r=1$, the performance degrades significantly, likely due to insufficient temporal information for accurate motion estimation. Increasing $r$ to 5 leads to a performance decline, though less severe than $r=1$. This degradation at $r=5$ can be attributed to the violation of the local linear motion assumption, which is fundamental to the CMax framework.
A.4.3 Computational Efficiency and Impact of Subinterval Number N
To investigate the impact of the number of subintervals N on our method, we conduct ablation studies on the Tanks and Temples dataset under the 2 FPS setting. The results are shown in Table 10 and Fig. 8. We test the speed on RTX2080ti. Note that in the t-PSNR figure in Fig 8, circular nodes represent pose-free methods, while diamond-shaped nodes indicate methods that require precomputed poses. The results in Table 10 reveal that increasing N improves PSNR, indicating that finer temporal resolution enhances reconstruction quality. However, this improvement comes at the cost of an extended training time, increasing from 1.3 hours at N=2 to 7 hours at N=6. Importantly, our method maintains real-time rendering performance (30+ FPS), matching CF-3DGS as both utilize the same efficient 3DGS rendering framework. This gives our approach a significant advantage over slower NeRF-based methods. Even at N=2, our approach outperforms baselines in both PSNR and rendering speed. Event-3DGS achieves a good balance of efficiency and performance, but it requires precomputed poses. The setting of N=3 offers a good balance, achieving higher PSNR than CF-3DGS with comparable training time. However, it’s important to note that the optimal N may vary for different frame rates and motion characteristics.
Table 10: Effect of the number of event subintervals N.
[…]
$N=2$ 23.56 30+ 1.3h
$N=3$ 24.43 30+ 3h
$N=6$ 24.81 30+ 7h
[…]
CF-3DGS 22.08 30+ 3h
Figure 8: Comparison of training time vs PSNR for various methods
Figure 9: Study of the different evalution metrics with respect to $\lambda_{cm}$
Figure 10: Study of the different evalution metrics with respect to $\lambda_{grad}$
Figure 11: Study of the different evalution metrics with respect to $\lambda_{PBA}$
Figure 12: Qualitative Results of the effectiveness of $\mathcal{L}_{EGM}$.
Figure 13: Qualitative Results of the effectiveness of $\mathcal{L}_{LEGM}$.
Figure 14: Qualitative Results of the effectiveness of $\mathcal{L}_{PBA}$.
A.4.4 Analysis of Loss Coefficients
Different components of our method play crucial roles in overall performance. To comprehensively understand these effects and determine optimal weight settings, we conducted a detailed analysis of the loss weights: $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$.
Contrast Maximization Coefficient $\lambda_{cm}$
As shown in Fig. 11, the contrast maximization coefficient $\lambda_{cm}$ significantly affects both NVS quality and pose estimation. We observe that a value of 0.1 achieves the best overall performance. Lower values (0.01) lead to decreased performance, likely due to insufficient utilization of motion information from events. Higher values (0.5) also result in performance degradation, possibly due to over-reliance on the contrast maximization term at the expense of other constraints. The optimal $\lambda_{cm}$ suggests that while event data is crucial, it should not dominate the reconstruction process. This balance allows our method to leverage the high temporal resolution of events without sacrificing the global supervision provided by frame data.
Gradient-based Loss Coefficient $\lambda_{grad}$
As shown in Fig. 11, the gradient-based loss coefficient $\lambda_{grad}$ shows optimal performance at 0.2. Lower values (0.1) slightly decrease performance, while higher values (0.3, 0.5) lead to more significant drops in both NVS quality and pose accuracy. This is likely because $\mathcal{L}_{EGM}$ and 3DGS primarily focus on rendering absolute brightness, and excessive gradient constraints may interfere with this process. Therefore, a moderate $\lambda_{grad}$ value is crucial to balance gradient information with the primary rendering objectives.
Photometric Bundle Adjustment Coefficient $\lambda_{PBA}$
As illustrated in Figure 11, our method demonstrates remarkable stability across $\lambda_{PBA}$ values ranging from 0.1 to 0.5. However, while $\mathcal{L}_{PBA}$ has minimal effect on PSNR and ATE metrics, its absence leads to noticeable color distortion artifacts in visual results. This underscores the importance of $\lambda_{PBA}$ in maintaining visual fidelity. Conversely, setting $\lambda_{PBA}$ to 1.0 results in significant performance degradation, indicating that overemphasis on cross-view consistency can be counterproductive.
A.5 Visualization of Ablation Studies
Due to space constraints, we provide comprehensive visualizations of our ablation studies in this section. Our visualizations demonstrate that while CF-3DGS can produce high-quality renderings at training views, it struggles with adjacent novel views. The rendered depth maps expose considerable inaccuracies in scene geometry reconstruction. With $\mathcal{L}_{EGM}$, as shown in Fig. 14, we observe marked improvements as it effectively harnesses continuous brightness change data captured by event cameras, compensating for inter-frame information loss. For $\mathcal{L}_{LEGM}$, as shown in Fig. 14, our visualizations show that without this component, the generated Image of Warped Events (IWE) exhibits blurring effects, indicating inaccurate pose estimation. This is particularly evident in self-similar regions with repetitive textures, which pose additional challenges. $\mathcal{L}_{LEGM}$ optimizes IWE sharpness, effectively improving both rendering quality and pose estimation. Regarding $\mathcal{L}_{PBA}$ and Fixed-GS, as shown in Fig. 14, our visualizations reveal that without these components, the reconstructed scene loses almost all color information. This occurs because event-guided optimization, with substantially more event frames than image frames, dominates scene reconstruction and suppresses color information. $\mathcal{L}_{PBA}$ alone, which reprojects neighboring video frames onto event frames, only partially addresses this imbalance. The Fixed-GS strategy separates color and structure optimization, and when combined with PBA, they complement each other to effectively resolve color distortion issues.
A.6 Additional Challenging Scenario Experiments
To further demonstrate event camera advantages, we constructed a FASTER scenario by selecting every 10th frame from the SLOW dataset and directly applied COLMAP pose estimation on these challenging sequences. We added COLMAP+3DGS as a comparison baseline. Notably, in the corner and outdoor scenes, several frames exhibit registration failures where the estimated poses greatly deviate from the overall video trajectory, indicating COLMAP’s limitations in challenging real-world scenarios.
Table 11: Rendering quality comparison across different speed scenarios on RealEv-DAVIS dataset.
Methods SLOW FAST FASTER
[…]
CF-3DGS 22.68 0.629 17.59 0.520 15.61 0.504
COLMAP+3DGS 23.27 0.633 19.79 0.548 17.25 0.521
EF-3DGS (Ours) 23.65 0.647 21.12 0.562 20.18 0.532
As shown in Table 11, the results demonstrate that: (1) COLMAP+3DGS rendering quality degrades significantly as motion speed increase confirming pose estimation inaccuracies in challenging conditions. (2) Compared to CF-3DGS, our method maintains substantially better rendering quality in high-speed scenarios, highlighting our method robustness and demonstrating the potential of event cameras in high-speed scenarios.
A.7 Necessity of Two-Stage Optimization
To validate the necessity of our two-stage Fixed-GS training strategy, we explore an alternative approach: incorporating a reference view rendering loss $\mathcal{L}_{\text{ref}}$ at each training iteration to address the color distortion challenge without requiring a separate optimization stage.
Given that events have much higher temporal resolution than RGB frames, the sparse color constraints from RGB images are largely overwhelmed by the abundant grayscale constraints from event data, ultimately resulting in color distortion. By incorporating $\mathcal{L}_{\text{ref}}$, the backpropagated gradients at each iteration would be augmented with color information from the reference view, potentially alleviating the color distortion issue. We conducted additional experiments comparing this approach with our two-stage strategy. As shown in Table 12, the results show that incorporating $\mathcal{L}_{\text{ref}}$ achieves comparable rendering quality. Although we cannot include visualizations here, our qualitative results confirm that the color distortion issue is effectively alleviated with this approach. However, this comes at a significant computational cost, increasing training time (3.8h vs 2.5h) due to additional reference view rendering at each iteration. Considering the modest performance difference relative to the substantial training overhead, we believe the two-stage optimization strategy offers a more efficient and practical solution.
Table 12: Comparison of different color correction strategies.
[…]
Baseline 23.09 0.70 0.38 2.7h
$\mathcal{L}_{event}+\mathcal{L}_{color}$ 24.11 0.73 0.36 3.8h
$\mathcal{L}_{event}$ + two-stage (Fixed-GS) strategy 23.96 0.72 0.36 2.5h
A.8 Additional Visualization
We present additional qualitative results on both Tanks and Temples and RealEv-DAVIS.
A.9 Limitations
Our method combines event streams with conventional video frames, which require the input data to be time-ordered. So it can not handle the unordered image input. And, our method relies on several key parameters (e.g., $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$) that may require manual tuning for optimal performance across different speed scenarios. This dependency on scene-specific parameter settings could limit the method’s adaptability to diverse environments. Future work should explore self-adaptive parameter adjustment strategies to enhance the method’s versatility and ease of use across various reconstruction tasks.
A.10 Broader Impacts
To the best of our knowledge, the proposed method will not have significant negative social impact. The proposed scene reconstruction method can be used to reconstruct and render some wild scenes. Users can use the video shot by their mobile phones as input to obtain an explicit 3D asset represented by a 3D Gaussian. This 3D asset can be used for subsequent editing, development, secondary creation for entertainment.
A.11 Data Availability
The datasets that support the findings of this study are available in the following repositories: Tanks and Temples tnt at https://www.tanksandtemples.org/ under CC BY 4.0 license, The code of choosen baseline, Nope-NeRF nopenerf is available at https://github.com/WU-CVGL/BAD-Gaussians under Apache-2.0 license, LocalRF localrf is available at https://localrf.github.io/ under Apache-2.0 license, Event-3DGS event3dgs is available at https://github.com/lanpokn/Event-3DGS under Apache-2.0 license, EvDeblurNeRF evdeblurnerf is available at https://github.com/uzh-rpg/EvDeblurNeRF under Apache-2.0 license, ENeRF enerf is available at https://github.com/knelk/enerf under MIT license, CF-3DGS cf3dgs is available at https://github.com/NVlabs/CF-3DGS under Apache-2.0 license, .
Figure 15: Qualitative comparison for novel view synthesis on Tanks and Temples. Our approach produces more realistic rendering results with fine-grained details. Better viewed when zoomed in.
Figure 16: Qualitative comparison for novel view synthesis on RealEv-DAVIS. Our approach produces more realistic rendering results with fine-grained details. Better viewed when zoomed in.
Figure 17: Pose Estimation Comparison on Tanks and Temples. We visualise the trajectory (3D plot) and relative rotation errors $\textrm{RPE}_{r}$ (bottom colour bar) of each method. We clip and normalize the $\textrm{RPE}_{r}$ by a quarter of the max $\textrm{RPE}_{r}$ across all results of each scene.
Figure 18: Pose Estimation Comparison on Tanks and Temples. We visualise the trajectory (3D plot) and relative rotation errors $\textrm{RPE}_{r}$ (bottom colour bar) of each method. We clip and normalize the $\textrm{RPE}_{r}$ by a quarter of the max $\textrm{RPE}_{r}$ across all results of each scene.
Figure 19: Pose Estimation Comparison on Tanks and Temples and RealEv-DAVIS. We visualise the trajectory (3D plot) and relative rotation errors $\textrm{RPE}_{r}$ (bottom colour bar) of each method. We clip and normalize the $\textrm{RPE}_{r}$ by a quarter of the max $\textrm{RPE}_{r}$ across all results of each scene.